# Weak Cosmic Censorship and Weak Gravity Conjectures in CFT Thermodynamics

### v1

@@ 5 CFT thermodynamics
$\begin{split}ds^{2}=\omega^{2}(-dt^{2}+\ell^{2}d\Omega_{d-2}^{2}).\end{split}$ (22)
@@
$\begin{split}\mathcal{V}=\Omega_{d-2}R^{d-2},\end{split}$ (23)
@@
$\begin{split}E=\frac{M}{\omega},\qquad\tilde{T}=\frac{T}{\omega},\qquad\tilde{S}=S,\qquad\tilde{Q}=\frac{Q\ell}{\sqrt{G}},\qquad\tilde{\Phi}=\frac{\Phi\sqrt{G}}{\omega\ell}.\end{split}$ (24)
@@
$\begin{split}\delta E=\tilde{\Phi}\delta\tilde{Q}+\mu\delta C-p\delta\mathcal{V}+\tilde{T}\delta S,\end{split}$ (25)
@@
$\begin{split}\mu=\frac{E-\tilde{T}S-\tilde{\Phi}\tilde{Q}}{C},\end{split}$ (26)
$\begin{split}p=\frac{E}{(d-2)\mathcal{V}}.\end{split}$ (27)
@@
$\Omega_{2}=4\pi,\qquad C=\frac{\ell^{2}}{4G},$ (29)
@@
$\begin{split}&\delta\bigg{(}\frac{M}{\omega}\bigg{)}=\frac{T}{\omega}\delta\bigg{(}\frac{A}{4G}\bigg{)}+\bigg{(}\frac{M}{\omega}-\frac{TS}{\omega}-\frac{Q\Phi}{\omega}-\frac{\Pi\gamma}{\omega}\bigg{)}\frac{\delta(\ell^{2}/4G)}{\ell^{2}/4G}\\ &-\frac{M}{2\omega}\frac{\delta(\Omega_{2}R^{2})}{(\Omega_{2}R^{2})}+\frac{\Phi\sqrt{G}}{\omega\ell}\delta\bigg{(}\frac{Q\ell}{\sqrt{G}}\bigg{)}+\frac{\Pi\sqrt{G}}{\omega\ell}\delta\bigg{(}\frac{\gamma\ell}{\sqrt{G}}\bigg{)}.\end{split}$ (30)
@@
$\begin{split}\delta E=\tilde{T}\delta\tilde{S}+\mu\delta C-p\delta\mathcal{V}+\tilde{\Phi}\delta\tilde{Q}+\tilde{\Pi}\delta\tilde{\gamma},\end{split}$ (31)
@@
$\begin{split}E=\frac{M}{\omega},\quad\tilde{S}=S,\quad\tilde{T}=\frac{T}{\omega},\quad\tilde{\Phi}=\frac{\Phi\sqrt{G}}{\omega\ell},\quad\tilde{Q}=\frac{Q\ell}{\sqrt{G}},\quad\tilde{\Pi}=\frac{\Pi\sqrt{G}}{\omega\ell},\quad\tilde{\gamma}=\frac{\gamma\ell}{\sqrt{G}},\end{split}$ (32)
@@
$\begin{split}E=\frac{4\pi CS+\pi^{2}\tilde{Q}^{2}+S^{2}}{2\pi\sqrt{CS\mathcal{V}}}+\frac{2\beta\sqrt{\pi C}}{\sqrt{\mathcal{V}}}\ln\left(\frac{\sqrt{S}}{\sqrt{\pi}\beta}\right),\end{split}$ (33)
$\tilde{T}=\bigg{(}\frac{\partial E}{\partial S}\bigg{)}_{\tilde{Q},C,\beta,\mathcal{V}}=\frac{4\pi CS-\pi^{2}\tilde{Q}^{2}+3S^{2}}{4\pi S^{3/2}\sqrt{C\mathcal{V}}}+\frac{\beta\sqrt{\pi C}}{S\sqrt{\mathcal{V}}},$ (34)
$\tilde{\Phi}=\bigg{(}\frac{\partial E}{\partial\tilde{Q}}\bigg{)}_{C,S,\beta,\mathcal{V}}=\frac{\pi\tilde{Q}}{\sqrt{CS\mathcal{V}}},$ (35)
$\mu=\bigg{(}\frac{\partial E}{\partial C}\bigg{)}_{\tilde{Q},S,\beta,\mathcal{V}}=\frac{4\pi CS-\pi^{2}\tilde{Q}^{2}-S^{2}}{4\pi C\sqrt{CS\mathcal{V}}}+\frac{\sqrt{\pi}\beta}{\sqrt{C\mathcal{V}}}\ln\bigg{(}\frac{\sqrt{S}}{\sqrt{\pi}\beta}\bigg{)},$ (36)
$\tilde{\Pi}=\bigg{(}\frac{\partial E}{\partial\tilde{\gamma}}\bigg{)}_{\tilde{Q},S,C,\mathcal{V}}=\frac{\beta\sqrt{\pi C}}{\tilde{\gamma}S\sqrt{\mathcal{V}}},$ (37)
$-p=\bigg{(}\frac{\partial E}{\partial\mathcal{V}}\bigg{)}_{\tilde{Q},S,C,\beta}=-\frac{E}{2\mathcal{V}}.$ (38)
@@
$\begin{split}\bigg{(}\frac{\partial\tilde{T}}{\partial S}\bigg{)}_{\tilde{Q},\mathcal{V},C,\beta}=0,\qquad\bigg{(}\frac{\partial^{2}\tilde{T}}{\partial S^{2}}\bigg{)}_{\tilde{Q},\mathcal{V},C,\beta}=0.\end{split}$ (39)
@@
$X:=(-4C^{3}+81\beta^{2}C^{2}+9\beta C^{2}\sqrt{81\beta^{2}-8C})^{\frac{1}{3}},\qquad S_{c}=\frac{1}{9}\left(\frac{2^{5/3}\pi C^{2}}{X}+4\pi C+2^{\frac{1}{3}}\pi X\right),$ (40)
$\tilde{Q}_{c}^{2}=\frac{-2^{10/3}C^{4}+2^{11/3}C^{3}X+24C^{2}X^{2}+2^{7/3}CX^{3}-2^{2/3}X^{4}}{81X^{2}}+\frac{8}{9}\beta C\sqrt{\frac{2^{5/3}C^{2}}{X}+4C+2^{\frac{1}{3}}X},$ (41)
$\tilde{T}_{c}=\frac{4\pi CS_{c}-\pi^{2}\tilde{Q}_{c}^{2}+3S_{c}^{2}}{4\pi S_{c}^{3/2}\sqrt{C\mathcal{V}}}+\frac{\beta\sqrt{\pi C}}{S_{c}\sqrt{\mathcal{V}}},\qquad E_{c}=\frac{4\pi CS_{c}+\pi^{2}\tilde{Q}_{c}^{2}+S_{c}^{2}}{2\pi\sqrt{CS_{c}\mathcal{V}}}+\frac{2\beta\sqrt{\pi C}}{\sqrt{\mathcal{V}}}\ln\bigg{(}\frac{\sqrt{S_{c}}}{\sqrt{\pi}\beta}\bigg{)}.$ (42)
@@
$F=E-\tilde{T}\tilde{S}=\frac{12\pi CS+\pi^{2}\tilde{Q}^{2}+5S^{2}}{4\pi\sqrt{CS\mathcal{V}}}+\frac{\beta\sqrt{\pi C}}{\sqrt{\mathcal{V}}}\bigg{[}-1+2\ln\bigg{(}\frac{\sqrt{S}}{\sqrt{\pi}\beta}\bigg{)}\bigg{]}.$ (43)
Figures 4 and 5 showcase the evolution of free energy as a function of temperature for diverse free parameter values of the RN AdS black hole, situated within the theoretical framework of PFDM in Conformal Field Theory. A notable swallowtail pattern is observed when $\tilde{Q}_{c}>\tilde{Q}=1.5$, signaling a first-order phase transition. Also, in $\tilde{Q}_{c}=\tilde{Q}$, we face with second phase transition. For enhanced visual clarity, we have separated the figures. Each figure is meticulously labeled with the relevant free parameters. By separating the figures, we aim to highlight and scrutinize the distinct features of each parameter set, thereby enabling a more granular analysis of the system’s dynamics under different conditions. The role of free parameters is pivotal in determining the physical attributes of the black hole.
@@
Figure 6 illustrates the Hawking temperature as a function of entropy for the RN AdS black hole in the context of PFDM within CFT. We set $C$ =$\mathcal{V}$=1. For Figure 6a, the parameters are chosen as $\beta=2$ and $\tilde{Q}_{c}=2.52$, whereas for Figure 6b, $\beta=1.5$ and $\tilde{Q}_{c}=2.15$ are selected. Each subplot in the figure discloses distinct zero points for varying free parameters across the entropy range $10<S<$ 1. A detailed analysis of the entropy in relation to the black hole’s free parameter values reveals a pronounced uniqueness in structural behavior for lower entropy values, denoted by $S$. This uniqueness is characterized by significant deviations in the thermal properties of the black hole. In contrast, as the entropy value $S$ ascends, the subplots exhibit a tendency to align, signifying a behavioral unification at higher entropy levels.
6 WGC &\&& WCCC
@@
The primary objective of this work is to challenge the WGC and WCCC in different spaces. So first we investigate the thermodynamic properties of Reissner-Nordström black holes in the presence PFDM, utilizing the principles of RPS and CFT thermodynamics. Then we explore the special ranges where these conjectures are compatible with each other in the $Q/M>\sqrt{G}$. We aim to explore the following key aspects: This comprehensive study aims to deepen our understanding of black hole physics in the presence of dark matter and energy, contributing to the broader field of quantum gravity and cosmology. This research objective is designed to be broad yet focused, allowing for a detailed exploration of the interconnected theories and their implications on the thermodynamics and stability of black holes in astrophysical scenarios influenced by dark matter.
Figure 7: (a) $\ell=0.5$,            (b) $\ell=1.5$,           (c) $\ell=2$
To substantiate the WGC within various phase spaces, we must establish certain specific relations. In the EPST, the condition is $H=\frac{Q^{2}}{M^{2}}-1>0$, and for the RPST and CFT, the criterion is $H=\frac{Q^{2}}{M^{2}}-G>0$. Consequently, we can pursue several approaches, among which thermodynamics and the critical values of black hole parameters are particularly significant. In light of this, we integrate the critical points delineated in prior sections into each respective space under the conditions mentioned above. Given that the critical values related with two parameters, namely $\gamma$ and $l$, it becomes evident that these parameters play pivotal roles in validating the WGC. As depicted in Figure 7, the ranges within which the WGC is satisfied—specifically ($H>0$) for EPST and RPST are the same. However, for CFT, this range is different and larger than other spaces.
@@
The WGC in CFT thermodynamic space is considered to yield a better result and is within a more acceptable range for several reasons and the coordination of the WGC and CFT can be checked in different works[110, 111, 112, 113]

### v2

@@ 4 CFT thermodynamics
$\begin{split}ds^{2}=\omega^{2}(-dt^{2}+\ell^{2}d\Omega_{d-2}^{2}).\end{split}$ (4.1)
@@
$\begin{split}\mathcal{V}=\Omega_{d-2}R^{d-2},\end{split}$ (4.2)
@@
$\begin{split}E=\frac{M}{\omega},\qquad\tilde{T}=\frac{T}{\omega},\qquad\tilde{S}=S,\qquad\tilde{Q}=\frac{Q\ell}{\sqrt{G}},\qquad\tilde{\Phi}=\frac{\Phi\sqrt{G}}{\omega\ell}.\end{split}$ (4.3)
@@
$\begin{split}\delta E=\tilde{\Phi}\delta\tilde{Q}+\mu\delta C-p\delta\mathcal{V}+\tilde{T}\delta S,\end{split}$ (4.4)
@@
$\begin{split}\mu=\frac{E-\tilde{T}S-\tilde{\Phi}\tilde{Q}}{C},\end{split}$ (4.5)
$\begin{split}p=\frac{E}{(d-2)\mathcal{V}}.\end{split}$ (4.6)
@@
$\Omega_{2}=4\pi,\qquad C=\frac{\ell^{2}}{4G},$ (4.8)
@@
$\begin{split}&\delta\bigg{(}\frac{M}{\omega}\bigg{)}=\frac{T}{\omega}\delta\bigg{(}\frac{A}{4G}\bigg{)}+\bigg{(}\frac{M}{\omega}-\frac{TS}{\omega}-\frac{Q\Phi}{\omega}-\frac{\Pi\gamma}{\omega}\bigg{)}\frac{\delta(\ell^{2}/4G)}{\ell^{2}/4G}\\ &-\frac{M}{2\omega}\frac{\delta(\Omega_{2}R^{2})}{(\Omega_{2}R^{2})}+\frac{\Phi\sqrt{G}}{\omega\ell}\delta\bigg{(}\frac{Q\ell}{\sqrt{G}}\bigg{)}+\frac{\Pi\sqrt{G}}{\omega\ell}\delta\bigg{(}\frac{\gamma\ell}{\sqrt{G}}\bigg{)}.\end{split}$ (4.9)
@@
$\begin{split}\delta E=\tilde{T}\delta\tilde{S}+\mu\delta C-p\delta\mathcal{V}+\tilde{\Phi}\delta\tilde{Q}+\tilde{\Pi}\delta\tilde{\gamma},\end{split}$ (4.10)
@@
$\begin{split}E=\frac{M}{\omega},\quad\tilde{S}=S,\quad\tilde{T}=\frac{T}{\omega},\quad\tilde{\Phi}=\frac{\Phi\sqrt{G}}{\omega\ell},\quad\tilde{Q}=\frac{Q\ell}{\sqrt{G}},\quad\tilde{\Pi}=\frac{\Pi\sqrt{G}}{\omega\ell},\quad\tilde{\gamma}=\frac{\gamma\ell}{\sqrt{G}},\end{split}$ (4.11)
@@
$\begin{split}E=\frac{4\pi CS+\pi^{2}\tilde{Q}^{2}+S^{2}}{2\pi\sqrt{CS\mathcal{V}}}+\frac{2\beta\sqrt{\pi C}}{\sqrt{\mathcal{V}}}\ln\left(\frac{\sqrt{S}}{\sqrt{\pi}\beta}\right),\end{split}$ (4.12)
$\tilde{T}=\bigg{(}\frac{\partial E}{\partial S}\bigg{)}_{\tilde{Q},C,\beta,\mathcal{V}}=\frac{4\pi CS-\pi^{2}\tilde{Q}^{2}+3S^{2}}{4\pi S^{3/2}\sqrt{C\mathcal{V}}}+\frac{\beta\sqrt{\pi C}}{S\sqrt{\mathcal{V}}},$ (4.13)
$\tilde{\Phi}=\bigg{(}\frac{\partial E}{\partial\tilde{Q}}\bigg{)}_{C,S,\beta,\mathcal{V}}=\frac{\pi\tilde{Q}}{\sqrt{CS\mathcal{V}}},$ (4.14)
$\mu=\bigg{(}\frac{\partial E}{\partial C}\bigg{)}_{\tilde{Q},S,\beta,\mathcal{V}}=\frac{4\pi CS-\pi^{2}\tilde{Q}^{2}-S^{2}}{4\pi C\sqrt{CS\mathcal{V}}}+\frac{\sqrt{\pi}\beta}{\sqrt{C\mathcal{V}}}\ln\bigg(\frac{\sqrt{S}}{\sqrt{\pi}\beta}\bigg{missing}),$ (4.15)
$\tilde{\Pi}=\bigg{(}\frac{\partial E}{\partial\tilde{\gamma}}\bigg{)}_{\tilde{Q},S,C,\mathcal{V}}=\frac{\beta\sqrt{\pi C}}{\tilde{\gamma}S\sqrt{\mathcal{V}}},$ (4.16)
$-p=\bigg{(}\frac{\partial E}{\partial\mathcal{V}}\bigg{)}_{\tilde{Q},S,C,\beta}=-\frac{E}{2\mathcal{V}}.$ (4.17)
@@
$\begin{split}\bigg{(}\frac{\partial\tilde{T}}{\partial S}\bigg{)}_{\tilde{Q},\mathcal{V},C,\beta}=0,\qquad\bigg{(}\frac{\partial^{2}\tilde{T}}{\partial S^{2}}\bigg{)}_{\tilde{Q},\mathcal{V},C,\beta}=0.\end{split}$ (4.18)
@@
$X:=(-4C^{3}+81\beta^{2}C^{2}+9\beta C^{2}\sqrt{81\beta^{2}-8C})^{\frac{1}{3}},\qquad S_{c}=\frac{1}{9}\left(\frac{2^{5/3}\pi C^{2}}{X}+4\pi C+2^{\frac{1}{3}}\pi X\right),$ (4.19)
$\tilde{Q}_{c}^{2}=\frac{-2^{10/3}C^{4}+2^{11/3}C^{3}X+24C^{2}X^{2}+2^{7/3}CX^{3}-2^{2/3}X^{4}}{81X^{2}}+\frac{8}{9}\beta C\sqrt{\frac{2^{5/3}C^{2}}{X}+4C+2^{\frac{1}{3}}X},$ (4.20)
$\tilde{T}_{c}=\frac{4\pi CS_{c}-\pi^{2}\tilde{Q}_{c}^{2}+3S_{c}^{2}}{4\pi S_{c}^{3/2}\sqrt{C\mathcal{V}}}+\frac{\beta\sqrt{\pi C}}{S_{c}\sqrt{\mathcal{V}}},\qquad E_{c}=\frac{4\pi CS_{c}+\pi^{2}\tilde{Q}_{c}^{2}+S_{c}^{2}}{2\pi\sqrt{CS_{c}\mathcal{V}}}+\frac{2\beta\sqrt{\pi C}}{\sqrt{\mathcal{V}}}\ln\bigg(\frac{\sqrt{S_{c}}}{\sqrt{\pi}\beta}\bigg{missing}).$ (4.21)
@@
$F=E-\tilde{T}\tilde{S}=\frac{12\pi CS+\pi^{2}\tilde{Q}^{2}+5S^{2}}{4\pi\sqrt{CS\mathcal{V}}}+\frac{\beta\sqrt{\pi C}}{\sqrt{\mathcal{V}}}\bigg{[}-1+2\ln\bigg(\frac{\sqrt{S}}{\sqrt{\pi}\beta}\bigg{missing})\bigg{]}.$ (4.22)
Figs. (4) and (5) showcase the evolution of free energy as a function of temperature for diverse free parameter values of the RN AdS black hole, situated within the theoretical framework of PFDM in Conformal Field Theory. A notable swallowtail pattern is observed when $\tilde{Q}_{c}>\tilde{Q}=1.5$, signaling a first-order phase transition. Also, in $\tilde{Q}_{c}=\tilde{Q}$, we face with second phase transition. For enhanced visual clarity, we have separated the figures. Each figure is meticulously labeled with the relevant free parameters. By separating the figures, we aim to highlight and scrutinize the distinct features of each parameter set, thereby enabling a more granular analysis of the system’s dynamics under different conditions. The role of free parameters is pivotal in determining the physical attributes of the black hole.
@@
Fig.(6) illustrates the Hawking temperature as a function of entropy for the RN AdS black hole in the context of PFDM within CFT. We set $C$ =$\mathcal{V}$=1. For Figure 6a, the parameters are chosen as $\beta=2$ and $\tilde{Q}_{c}=2.52$, whereas for Fig. (LABEL:6b), $\beta=1.5$ and $\tilde{Q}_{c}=2.15$ are selected. Each subplot in the figure discloses distinct zero points for varying free parameters across the entropy range $10<S<$ 1. A detailed analysis of the entropy of the black hole’s free parameter values reveals a pronounced uniqueness in structural behavior for lower entropy values, denoted by $S$. This uniqueness is characterized by significant deviations in the thermal properties of the black hole. In contrast, as the entropy value $S$ ascends, the subplots tend to align, signifying a behavioral unification at higher entropy levels. The potential violation of the second law of thermodynamics in the context of black hole thermodynamics has been a topic of significant interest, particularly in the various frameworks. In EPST, the cosmological constant is treated as a thermodynamic variable analogous to pressure, and its conjugate variable is the volume. This framework allows for a richer thermodynamic description of black holes, including phase transitions similar to those in conventional thermodynamics. RPST is a newer framework that excludes the $PdV$ term from the first law of black hole thermodynamics, focusing instead on other thermodynamic variables such as the central charge and its conjugate chemical potential. In the context of the AdS/CFT correspondence, black hole thermodynamics can be studied using the dual CFT. This approach provides a powerful tool for understanding the microscopic origins of black hole entropy and the second law. in[128, 129, 130] researchers have observed violations of the second law under certain conditions, particularly for extremal and near-extremal black holes. These violations are often linked to the behavior of the black hole’s entropy and the thermodynamic volume. More recent work[131, 132], suggests that these violations can be resolved by considering additional thermodynamic variables or by imposing certain constraints on the system. For instance, the inclusion of quantum corrections or the consideration of higher-dimensional spacetimes can restore the validity of the second law. Also shows potential violations of the second law, particularly in the context of charged black holes. These violations are often associated with the specific thermodynamic paths taken by the system. The researchers indicate that by carefully choosing the thermodynamic variables and paths, the second law can still hold. This involves ensuring that the entropy production is always non-negative, even in the presence of complex interactions and phase transitions. Recent works suggest that by considering the full spectrum of states in the CFT and including quantum corrections, the second law can be upheld. This involves ensuring that the entropy of the black hole and the dual CFT system always increases, even under extreme conditions. The potential violation of the second law of thermodynamics in black hole systems is a complex issue that depends heavily on the chosen thermodynamic framework. While violations have been observed under certain conditions, recent research indicates that these can often be resolved by considering additional variables, quantum corrections, or specific constraints. The frameworks of EPST, RPST, and CFT phase spaces each offer unique insights and mechanisms for ensuring the validity of the second law. Therefore, we consider an additional variable for the black hole, referred to as PFDM. This variable not only resolves the conflict between the WGC and the WCCC but also addresses this violation.
5 WGC and WCCC
@@
The primary objective of this work is to challenge the WGC and WCCC in different spaces. So first we investigate the thermodynamic properties of Reissner-Nordström black holes in the presence PFDM, utilizing the principles of RPS and CFT thermodynamics. Then we explore the special ranges where these conjectures are compatible with each other in the $Q/M>\sqrt{G}$. Our goal is to delve into some crucial concepts. This research is designed to enhance our comprehension of black hole physics, particularly in the context of dark matter and dark energy. We also study quantum gravity viz relation between quantum mechanics and cosmology through WGC and WCCC. So, it allows for an in-depth examination of the interlinked theories and their impact on the thermodynamics and stability of black holes.
Figure 7: Compatibility range of the weak gravity conjecture with different thermodynamics (EPST, RPST, CFT) according to free parameters. (a) $\ell=0.5$,            (b) $\ell=1.5$,           (c) $\ell=2$
[…]
@@
As can be seen above the WGC in CFT thermodynamic space yields a better result and is within a more acceptable range. The coordination of the WGC and CFT can be checked in different works[125, 126, 113, 127].
The presence of PFDM plays a pivotal role in reconciling the WGC and the WCCC within the context of RN black holes. We analyze how PFDM helps resolve the conflict between these two conjectures.
WGC posits that gravity should be the weakest force in any consistent theory of quantum gravity. It implies that there must be a particle with a charge-to-mass ratio greater than or equal to that of an extremal black hole, ensuring that black holes can decay. Also WCCC asserts that singularities arising from gravitational collapse are hidden within event horizons, preventing them from being observed from the rest of spacetime. This ensures the predictability of physical laws. PFDM modifies the spacetime geometry around RN black holes, which in turn affects the conditions under which the WGC and WCCC are satisfied. In the absence of PFDM, an RN black hole can either have two event horizons (if $Q^{2}/M^{2}\leq 1$) or none (if $Q^{2}/M^{2}>1$). The latter scenario results in a naked singularity, violating the WCCC. When PFDM is present, it alters the metric of the black hole, ensuring that the singularity is always covered by an event horizon, thus satisfying the WCCC. This is achieved by introducing a parameter $\gamma$ which represents the influence of PFDM.
PFDM changes the conditions for the existence of event horizons. By adjusting the parameter $\gamma$, the RN black hole can maintain its event horizons even when $Q^{2}/M^{2}>1$, preventing the formation of a naked singularity and thus upholding the WCCC. There exists a critical value of $\gamma$, denoted as $\gamma_{ext}$, which makes the RN black hole extremal. At this value, the black hole has an event horizon, ensuring that both the WGC and WCCC are satisfied. The parameter $\gamma$ and the radius of AdS space $l$ have a linear relationship with the range compatibility of the WGC. This means that the presence of PFDM expands the range of parameters for which the WGC is satisfied, making it more consistent with the WCCC. PFDM effectively reconciles the WGC and WCCC by modifying the spacetime geometry around RN black holes, ensuring that singularities are always hidden within event horizons and that gravity remains the weakest force. This dual satisfaction of the conjectures demonstrates the robustness of these theoretical frameworks when PFDM is considered.